OpenStudy (anonymous):

trying to determine the proper way to answer this question: Find ________ - /36 / 25 of course the brackets aren't suppose to be like the dashes I had to use to post the question. lol

OpenStudy (anonymous):

is it -\(\sqrt{36/25}\)..?

geerky42 (geerky42):

HINT: \(\large 36 = 6^2, ~ 25 = 5^2\)

OpenStudy (anonymous):

yes it is I understand the answers but are they written in a bracket? is it one bracket or two?

OpenStudy (anonymous):



OpenStudy (anonymous):

its simply 6/5..

OpenStudy (anonymous):

-6/5 rather..

geerky42 (geerky42):

\[\sqrt{\dfrac{36}{25}} = \sqrt{\dfrac{6^2}{5^2}} = \sqrt{\left(\dfrac{6}{5}\right)^2} = \boxed{\dfrac{6}{5}}\]

geerky42 (geerky42):

assume there is no negative sign.

OpenStudy (anonymous):

there is a negative sign outside of the first bracket

OpenStudy (anonymous):

\[-\sqrt{36/25}\]

geerky42 (geerky42):

so then asnwer is \(-\dfrac{6}{5}\) Is this what you are looking for?

OpenStudy (anonymous):

ok great. I just wasn't sure if it needed the other brackets thank you